# MeshSplat: Generalizable Sparse-View Surface Reconstruction via Gaussian Splatting

### v1

@@ Abstract
Figure 1: Given sparse-view images as input, MeshSplat can directly predict the scene geometry and efficiently extract the scene mesh. Compared to MVSplat [10] and other state-of-the-art methods, Meshplat achieves more consistent and precise mesh extraction in generalizable sparse-view surface reconstruction.
@@ 1 Introduction
Surface reconstruction of 3D scenes is a fundamental task in 3D vision, with a wide range of applications in downstream tasks such as AR/VR and embodied AI. Recently, based on NeRF [35] and 3DGS [26], numerous per-scene optimized surface reconstruction methods [52, 21, 6, 67, 19] have been proposed. However, these methods struggle to robustly reconstruct scenes with only sparse views as input. This is because sparse views can only provide limited multi-view geometric constraints, which are insufficient to perform high-quality per-scene geometry optimization. Therefore, it is necessary to construct a feed-forward pipeline to learn geometric priors from diverse scenarios for generalizable sparse-view surface reconstruction.
To address this issue, several meaningful attempts have been made, which are mainly NeuS-based methods [52]. For instance, SparseNeuS [32] constructs geometric volumes to estimate implicit SDF fields, which are then transformed to mesh. However, due to the inefficiency of implicit SDF representation and neural rendering, NeuS-based methods are limited to object-centric scenes and suffer from long rendering times. Compared with NeuS, Gaussian splatting can achieve not only faster rendering speeds but also better rendering quality. However, the vanilla Gaussian splatting framework falls short in representing surface geometry [21]. Therefore, existing Gaussian splatting-based methods [3, 10, 56] mainly focus on generalizable novel view synthesis, while methods applying Gaussian splatting to generalizable sparse-view surface reconstruction remain unexplored.
To realize high-quality surface reconstruction, the network requires not only realistic appearance rendering but also precise geometry recovery. More importantly, the former task usually has more available training data, while the latter requires expensive geometric labeling. This raises a question: is it possible to learn generalizable geometric priors from novel view synthesis task? Motivated by recent work 2DGS [21], we find an effective way to address this issue: using 2DGS as a bridge between novel view synthesis (NVS) and surface reconstruction. Compared with 3DGS, 2DGS is naturally more suitable for representing thin surfaces and can be easily used to extract surface mesh beyond rendering novel views, as shown in Figure 2 (a) and (b). To this end, we decide to construct a feed-forward framework to predict pixel-aligned 2DGS for generalizable sparse-view surface reconstruction.
@@ 2 Related Work
Early 3D scene methods [23, 18, 37] employ signed distance functions or occupancy fields to represent 3D scenes. NeRF [35] utilizes neural networks to construct the 3D radiance field for scene representation and renders RGB images by volume rendering, which inspires many subsequent works in 3D vision [41, 2, 40, 36, 20, 29, 47]. However, NeRF-based methods suffer from long-time inefficient rendering. 3DGS [26] models 3D scenes by explicit ellipsoidal Gaussian primitives and introduces a differentiable rendering method based on alpha-blending, achieving significant breakthroughs in novel view synthesis tasks. It has become the core 3D scene representation for many tasks, such as novel view synthesis [43, 16, 33], 3D assets generation [49, 65], 3D scene understanding [75, 64] and dynamic scene reconstruction [22, 57, 34, 61, 74]. Despite their success, original NeRF and 3DGS require per-scene optimization, lacking cross-scene generalization capabilities.
@@ 2.2 Generalizable 3D Reconstruction
Extensive works have proposed generalizable 3D reconstruction frameworks by training neural networks on large-scale datasets to learn cross-scene priors. Early approaches [4, 53, 66, 9, 58] primarily combined feed-forward neural networks with Neural Radiance Fields (NeRF). However, these methods often suffer from the inefficiencies inherent in NeRF training and rendering, which limit their performance. More recently, methods such as DUSt3R and its subsequent extensions [54, 13, 60, 51, 31] leverage Transformer-based architectures trained on large-scale 3D-annotated datasets to predict 3D point maps from input 2D images. While effective in generating point-based 3D representations, these approaches cannot generalize to novel views and are not suitable for surface reconstruction tasks. Methods like LVSM [24] support novel view synthesis (NVS) tasks, but they are limited to generating RGB images and thus cannot be directly applied to surface reconstruction. With the introduction of 3D Gaussian Splatting (3DGS), a new class of methods [3, 48, 10, 30, 55, 70], where MVSplat [10] represents a prominent approach, employ 3DGS as the 3D scene representation for generalizable 3D reconstruction. They usually first predict depth maps from input views, and then use back-projected 3D points as centers to construct pixel-aligned 3D Gaussian representations. However, despite their promising performance in NVS tasks, these generalizable frameworks remain inadequate for mesh extraction and surface reconstruction tasks.
@@ 2.3 Surface Reconstruction
Traditional surface reconstruction methods [25, 17] primarily model 3D scene geometry through explicit feature matching. With the rise of neural networks, many works [52, 42, 27, 62, 37, 38, 63] reconstruct scene surface by estimating neural occupancy fields or signed distance functions (SDF) by neural networks, and render novel views through surface rendering or volume rendering techniques. Recently, many works have extended Gaussian Splatting to mesh extraction. These methods can be categorized into two types: methods that directly use ellipsoidal 3DGS to extract meshes [19, 67, 68, 69], and methods that use flattened 3DGS or 2DGS [21, 6, 50, 5, 12, 7, 15]. However, these methods fail to generate high-quality meshes under sparse viewpoint inputs and lack cross-scene generalization. SparseNeuS [32] and subsequent works [39, 59, 44, 8, 28], inspired by generalizable novel view synthesis approaches, achieve generalizable surface reconstruction based on NeuS [52]. These methods extract feature maps from input images to construct a 3D geometry volume, which is used to obtain neural SDF fields and output the mesh of the scene. However, these methods exhibit poor efficiency in mesh extraction and are limited to object-centric scenes. 2DGS [21] outperforms NeuS-based methods in mesh extraction tasks, but how to generalize 2DGS to different scenes with sparse inputs still remains unexplored. Therefore, we propose MeshSplat, an end-to-end generalizable sparse-view surface reconstruction framework based on 2DGS.
@@ 3.2 Cost Volume Construction
Following MVSplat [10], we first use a CNN to extract a feature map from each input view. Then, we employ a Transformer that incorporates multi-view cross-attention to enable information exchanges across different views. As a result, we obtain the feature map $\{F_{i}\}_{i=1}^{2}$ for each view with a size of $\frac{H}{4}\times\frac{W}{4}\times C$.
@@ 3.3 Gaussian Prediction Network
Next, we discuss how to derive the attributes of pixel-aligned 2DGS from cost volumes. Following MVSplat [10], we first use U-Net with attention layers to refine the cost volume, taking feature maps as additional inputs. Using the refined cost volume, we can compute the final depth maps using equation (3), and then perform back-projection to obtain the 3D positions of the pixel-aligned 2DGS.
In addition, 2DGS requires more precise orientation prediction than the ellipsoid-like 3DGS. Based on this, we introduce an additional normal prediction network to predict the normal maps of the input images, which are further converted to the 2DGS rotation quaternions. We use a lightweight CNN $\phi_{\mathrm{rot}}$ to predict the orientations of the pixel-aligned 2DGS. The network takes the refined cost volumes, feature maps, and original RGB images at three different scales as input, and processes them through convolutional layers with residual connections. Finally the rotation quaternions $q$ for all three scales are obtained, which are further converted into normal vectors $n$ for each scale, along with kappa $\kappa$ which reflect the uncertainty of predicted normal vectors following [1]:
@@
For the remaining attributes of 2DGS, following [10], we predict them through a two-layer convolutional gaussian head, taking the refined cost volumes, feature maps, and original RGB images as input.
@@ 3.5 Uncertainty-Guided NLL Normal Loss
Due to the lack of regularization for normals, the predicted orientations of 2DGS still suffer from degradation, preventing the 2DGS from aligning precisely with the actual surfaces. To address this issue, we use the uncertainty-guided normal negative log-likelihood (NLL) loss proposed by [1] to supervise predicted normal maps at each scale, which is defined as:
@@
where $n_{i}$, $\kappa_{i}$, $\hat{n}_{i}$ stands for the predicted normal vector, kappa and pseudo ground-truth normal vector of pixel $i$. Following [1], we sample pixels based on their $\kappa_{i}$, forming the sampled sets $P_{\mathrm{sample}}$, and only apply NLL loss on $P_{\mathrm{sample}}$. We use an off-the-shelf monocular normal estimation model Omnidata [14] to provide pseudo ground-truth normal maps as supervision. Finally, the normal loss $\mathcal{L}_{\mathrm{normal}}$ is computed as the average of NLL losses in sampled points. Details can be found in supplementary materials.
@@ 3.6 Training and Inference
During inference, 2DGS is first reconstructed through our network taken two images as input. Subsequently, we follow the methods of 2DGS original paper [21] to extract the scene mesh.
@@
SparseNeuS [32] 5.9736 0.0141 0.1146 0.0248 11.93 0.0001 0.0145 0.0002
MVSNeRF [4] 0.6139 0.1390 0.1548 0.1407 0.5761 0.1320 0.2053 0.1514
PixelSplat [3] 1.4423 0.1067 0.0903 0.0944 0.3285 0.2599 0.3597 0.2948
MVSplat** [10] 0.4038 0.3986 0.2633 0.3139 - - - -
MVSplat [10] 0.4015 0.3949 0.2607 0.3100 0.3748 0.1992 0.2282 0.2095
@@
SparseNeuS [32] 13.16 0.0001 0.0112 0.0001 15.21 0.0001 0.0104 0.0002
MVSplat [10] 0.4506 0.1333 0.1591 0.1418 1.089 0.0528 0.0695 0.0564
@@ 4.1 Settings
Datasets. We train and evaluate our model on Re10K [73] and Scannet [11] datasets. Re10K is a large-scale home walkthrough multi-view dataset obtained from YouTube, containing 67,477 training scenes and 7,289 testing scenes. Scannet is a multi-view real-world indoor dataset, and following [71], we use 100 scenes for training and 8 scenes for testing. To evaluate the model generalizability, we also perform cross-dataset evaluations from Re10K to Scannet and Replica [46], a multi-view synthetic indoor dataset. Following [72], we use 8 scenes for testing in Replica dataset.
@@ 4.2 Main Results
Mesh Reconstruction. To enable comparisons with the state-of-the-arts, we transfer four methods from generalizable novel view synthesis to mesh reconstruction as our comparison. For fair comparison, we also apply TSDF fusion to extract meshes for the compared methods. Table 1 presents the experimental results on the two datasets. The experiments demonstrate that our method significantly outperforms these methods in all metrics, indicating that our approach has strong capabilities in surface reconstruction with sparse-view input. Figures 5 and 5 show the visualizations on Re10K and Scannet. For the compared methods, their extracted meshes suffer from holes and uneven surfaces. In contrast, MeshSplat can accurately reconstruct the meshes of complex scenes with smoother surfaces, demonstrating that our model can capture scene structure well even with sparse-view inputs. More qualitative comparisons can be found in Appendix.
Cross-Dataset Generalization. To validate the generalization ability, we conduct zero-shot transfer experiments on the Scannet and Replica datasets respectively, using the model trained on the Re10K dataset. As shown in Table 2 and Fig 6, our method provides significant improvements compared to the baselines, demonstrating the satisfactory generalizability of MeshSplat across diverse scenarios.
@@ 4.3 Ablations
Table 3: Results on Depth and Normal Prediction Experiments. It is clear that MeshSplat can predict depth and normal maps of input views correctly.
@@
Main Components. Table 4 and Figure 8 (a) present the results of our ablation studies conducted on the Scannet dataset. The first two lines of the table indicate that 2DGS is more suitable for mesh reconstruction tasks compared to 3DGS. Moreover, by incorporating our proposed WCD loss and normal prediction network, 2DGS can align better with the actual surfaces. These enhancements significantly improve surface fidelity and reduce geometric artifacts in the final output. Overall, these results demonstrate that each proposed component plays a complementary role in boosting the reconstruction quality.
Gaussian normals. To further demonstrate the necessities of the normal prediction network, we present comparisons on normal maps rendered by Gaussian Splatting in Figure 8 (b). Following [6], we define the normal vector of 3DGS as the direction of the axis which has the minimum value of scales. 2DGS suffer from more significant degradation in normals compared to 3DGS when no regularization is applied. After incorporating the normal prediction network and its corresponding regularization loss, the predicted orientations of 2DGS become more accurate.
Figure 8: Qualitative Ablation Studies. (a) Ablation studies on our proposed modules. (b) Comparisons of normal maps rendered by 2DGS and 3DGS.

### v2

@@ Abstract
Figure 1: Given sparse-view images as input, MeshSplat can directly predict the scene geometry and efficiently extract the scene mesh. Compared to MVSplat [chen2024mvsplat] and other state-of-the-art methods, Meshplat achieves more consistent and precise mesh extraction in generalizable sparse-view surface reconstruction.
@@ 1 Introduction
Surface reconstruction of 3D scenes is a fundamental task in 3D vision, with a wide range of applications in downstream tasks such as AR/VR and embodied AI. Recently, based on NeRF [mildenhall2021nerf] and 3DGS [kerbl20233d], numerous per-scene optimized surface reconstruction methods [wang2021neus, huang20242d, chen2023neusg, yu2024gsdf, guedon2024sugar] have been proposed. However, these methods struggle to robustly reconstruct scenes with only sparse views as input. This is because sparse views can only provide limited multi-view geometric constraints, which are insufficient to perform high-quality per-scene geometry optimization. Therefore, it is necessary to construct a feed-forward pipeline to learn geometric priors from diverse scenarios for generalizable sparse-view surface reconstruction.
To address this issue, several meaningful attempts have been made, which are mainly NeuS-based methods [wang2021neus]. For instance, SparseNeuS [long2022sparseneus] constructs geometric volumes to estimate implicit SDF fields, which are then transformed to mesh. However, due to the inefficiency of implicit SDF representation and neural rendering, NeuS-based methods are limited to object-centric scenes and suffer from long rendering times. Compared with NeuS, Gaussian splatting can achieve not only faster rendering speeds but also better rendering quality. However, the vanilla Gaussian splatting framework falls short in representing surface geometry [huang20242d]. Therefore, existing Gaussian splatting-based methods [charatan2024pixelsplat, chen2024mvsplat, wewer2024latentsplat] mainly focus on generalizable novel view synthesis, while methods applying Gaussian splatting to generalizable sparse-view surface reconstruction remain unexplored.
To realize high-quality surface reconstruction, the network requires not only realistic appearance rendering but also precise geometry recovery. More importantly, the former task usually has more available training data, while the latter requires expensive geometric labeling. This raises a question: is it possible to learn generalizable geometric priors from novel view synthesis task? Motivated by recent work 2DGS [huang20242d], we find an effective way to address this issue: using 2DGS as a bridge between novel view synthesis (NVS) and surface reconstruction. Compared with 3DGS, 2DGS is naturally more suitable for representing thin surfaces and can be easily used to extract surface mesh beyond rendering novel views, as shown in Figure 2 (a) and (b). To this end, we decide to construct a feed-forward framework to predict pixel-aligned 2DGS for generalizable sparse-view surface reconstruction.
@@ 2 Related Work
Early 3D scene methods [jiang2020local, genova2020local, cheng2025bridge] employ point clouds, signed distance functions or occupancy fields to represent 3D scenes. NeRF [mildenhall2021nerf] utilizes neural networks to construct the 3D radiance field for scene representation and renders RGB images by volume rendering, which inspires many subsequent works in 3D vision [barron2021mip, muller2022instant, honglrm]. However, NeRF-based methods suffer from long-time inefficient rendering. 3DGS [kerbl20233d] models 3D scenes by explicit ellipsoidal Gaussian primitives and introduces a differentiable rendering method based on alpha-blending. It has become the core 3D scene representation for many tasks, such as novel view synthesis [ren2025octree, lu2024scaffold], 3D assets generation [tangdreamgaussian] and dynamic scene reconstruction [huang2024sc, zhu2024motiongs]. Despite their success, original NeRF and 3DGS require per-scene optimization, lacking cross-scene generalization capabilities.
@@ 2.2 Generalizable 3D Reconstruction
Extensive works have proposed generalizable 3D reconstruction frameworks by training neural networks on large-scale datasets to learn cross-scene priors. Early approaches [chen2021mvsnerf, wang2021ibrnet, xu2024murf] primarily combined feed-forward neural networks with Neural Radiance Fields (NeRF). However, these methods often suffer from the inefficiencies inherent in NeRF training and rendering, which limit their performance. More recently, methods such as DUSt3R and its subsequent extensions [wang2024dust3r, duisterhof2025mast3r, wang2025vggt] leverage Transformer-based architectures trained on large-scale 3D-annotated datasets to predict 3D point maps from input 2D images. While effective in generating point-based 3D representations, these approaches cannot generalize to novel views and are not suitable for surface reconstruction tasks. Methods like LVSM [jinlvsm] support novel view synthesis (NVS) tasks, but they are limited to generating RGB images and thus cannot be directly applied to surface reconstruction. With the introduction of 3D Gaussian Splatting (3DGS), a new class of methods [charatan2024pixelsplat, chen2024mvsplat, wang2024freesplat, wang2025zpressor], where MVSplat [chen2024mvsplat] represents a prominent approach, employ 3DGS as the 3D scene representation for generalizable 3D reconstruction. They usually first predict depth maps from input views, and then use back-projected 3D points as centers to construct pixel-aligned 3D Gaussian representations. PM loss [shi2025revisiting] further regularizes the predicted 3DGS positions by introducing VGGT priors. However, despite their promising performance in NVS tasks, these generalizable frameworks remain inadequate for mesh extraction and surface reconstruction tasks.
@@ 2.3 Surface Reconstruction
Traditional surface reconstruction methods [kazhdan2013screened, furukawa2009accurate] primarily model 3D scene geometry through explicit feature matching. With the rise of neural networks, many works [wang2021neus, reiser2024binary, li2023neuralangelo, yariv2021volume] reconstruct scene surface by estimating neural occupancy fields or signed distance functions (SDF) by neural networks, and render novel views through surface rendering or volume rendering techniques. Recently, many works have extended Gaussian Splatting to mesh extraction. These methods can be categorized into two types: methods that directly use ellipsoidal 3DGS to extract meshes [guedon2024sugar, yu2024gsdf, yu2024gaussian, zhu2025objectgs], and methods that use flattened 3DGS or 2DGS [huang20242d, chen2023neusg, chen2024pgsr]. However, these methods fail to generate high-quality meshes under sparse viewpoint inputs and lack cross-scene generalization. SparseNeuS [long2022sparseneus] and subsequent works [peng2023gens, xu2023c2f2neus, chen2024neuralrecon], inspired by generalizable novel view synthesis approaches, achieve generalizable surface reconstruction based on NeuS [wang2021neus]. These methods extract feature maps from input images to construct a 3D geometry volume, which is used to obtain neural SDF fields and output the mesh of the scene. However, these methods exhibit poor efficiency in mesh extraction and are limited to object-centric scenes. 2DGS [huang20242d] outperforms NeuS-based methods in mesh extraction tasks, but how to generalize 2DGS to different scenes with sparse inputs still remains unexplored. Therefore, we propose MeshSplat, an end-to-end generalizable sparse-view surface reconstruction framework based on 2DGS.
@@ 3.2 Cost Volume Construction
Following MVSplat [chen2024mvsplat], we first use a CNN to extract a feature map from each input view. Then, we employ a Transformer that incorporates multi-view cross-attention to enable information exchanges across different views. As a result, we obtain the feature map $\{F_{i}\}_{i=1}^{2}$ for each view with a size of $\frac{H}{4}\times\frac{W}{4}\times C$.
@@ 3.3 Gaussian Prediction Network
Next, we discuss how to derive the attributes of pixel-aligned 2DGS from cost volumes. Following MVSplat [chen2024mvsplat], we first use U-Net with attention layers to refine the cost volume, taking feature maps as additional inputs. Using the refined cost volume, we can compute the final depth maps using equation (3), and then perform back-projection to obtain the 3D positions of the pixel-aligned 2DGS.
In addition, 2DGS requires more precise orientation prediction than the ellipsoid-like 3DGS. Based on this, we introduce an additional normal prediction network to predict the normal maps of the input images, which are further converted to the 2DGS rotation quaternions. We use a lightweight CNN $\phi_{\mathrm{rot}}$ to predict the orientations of the pixel-aligned 2DGS. The network takes the refined cost volumes, feature maps, and original RGB images at three different scales as input, and processes them through convolutional layers with residual connections. Finally the rotation quaternions $q$ for all three scales are obtained, which are further converted into normal vectors $n$ for each scale, along with kappa $\kappa$ which reflect the uncertainty of predicted normal vectors following [bae2021estimating]:
@@
For the remaining attributes of 2DGS, following [chen2024mvsplat], we predict them through a two-layer convolutional gaussian head, taking the refined cost volumes, feature maps, and original RGB images as input.
@@ 3.5 Uncertainty-Guided NLL Normal Loss
Due to the lack of regularization for normals, the predicted orientations of 2DGS still suffer from degradation, preventing the 2DGS from aligning precisely with the actual surfaces. To address this issue, we use the uncertainty-guided normal negative log-likelihood (NLL) loss proposed by [bae2021estimating] to supervise predicted normal maps at each scale, which is defined as:
@@
where $n_{i}$, $\kappa_{i}$, $\hat{n}_{i}$ stands for the predicted normal vector, kappa and pseudo ground-truth normal vector of pixel $i$. Following [bae2021estimating], we sample pixels based on their $\kappa_{i}$, forming the sampled sets $P_{\mathrm{sample}}$, and only apply NLL loss on $P_{\mathrm{sample}}$. We use an off-the-shelf monocular normal estimation model Omnidata [eftekhar2021omnidata] to provide pseudo ground-truth normal maps as supervision. Finally, the normal loss $\mathcal{L}_{\mathrm{normal}}$ is computed as the average of NLL losses in sampled points. Details can be found in the Appendix.
@@ 3.6 Training and Inference
During inference, 2DGS is first reconstructed through our network. Then we follow 2DGS original paper [huang20242d] to extract the scene mesh.
@@
SparseNeuS [long2022sparseneus] 6.0473 0.0012 0.0097 0.0020 7.1860 0.0056 0.1964 0.0107
MVSNeRF [chen2021mvsnerf] 0.6139 0.1390 0.1548 0.1407 0.5761 0.1320 0.2053 0.1514
pixelSplat [charatan2024pixelsplat] 1.4423 0.1067 0.0903 0.0944 0.3285 0.2599 0.3597 0.2948
MVSplat** [chen2024mvsplat] 0.4038 0.3986 0.2633 0.3139 - - - -
MVSplat [chen2024mvsplat] 0.4015 0.3949 0.2607 0.3100 0.3748 0.1992 0.2282 0.2095
@@
SparseNeuS [long2022sparseneus] 8.3055 0.0003 0.0166 0.0006 12.72 0.0002 0.0099 0.0003
MVSplat [chen2024mvsplat] 0.4506 0.1333 0.1591 0.1418 1.089 0.0528 0.0695 0.0564
@@ 4.1 Settings
Datasets. We train and evaluate our model on Re10K [zhou2018stereo] and Scannet [dai2017scannet] datasets. Re10K is a large-scale home walkthrough multi-view dataset obtained from YouTube, containing 67,477 training scenes and 7,289 testing scenes. Scannet is a multi-view real-world indoor dataset, and following [zhang2022nerfusion], we use 100 scenes for training and 8 scenes for testing. To evaluate the model generalizability, we also perform cross-dataset evaluations from Re10K to Scannet and Replica [straub2019replica], a multi-view synthetic indoor dataset. Following [zhi2021place], we use 8 scenes for testing in Replica dataset.
@@ 4.2 Main Results
Mesh Reconstruction. To enable comparisons with the state-of-the-arts, we transfer four methods from generalizable novel view synthesis to mesh reconstruction as our comparison. For fair comparison, we also apply TSDF fusion for them except for SparseNeuS as it can extract meshes directly. Table 1 presents the experimental results on the two datasets. The experiments demonstrate that MeshSplat significantly outperforms these methods in all metrics, indicating that our approach has strong capabilities in surface reconstruction with sparse-view input.
Figures 5 and 5 show the visualizations on Re10K and Scannet. For the compared methods, their extracted meshes suffer from holes and uneven surfaces. In contrast, MeshSplat can accurately reconstruct the meshes of complex scenes with smoother surfaces, demonstrating that our model can capture scene structure well even with sparse-view inputs. More comparisons can be found in Appendix.
Cross-Dataset Generalization. To validate the generalization ability, We also conduct zero-shot transfer experiments on Scannet and Replica datasets respectively, using the model trained on the Re10K dataset. As shown in Table 2 and Figure 6, our method still provides significant improvements compared to the baselines, demonstrating the satisfactory generalizability of MeshSplat across diverse scenarios.
@@ 4.3 Ablations
Table 3: Results on Depth and Normal Prediction. MeshSplat can predict depth and normal maps correctly.
@@
Figure 8: Qualitative Ablation Studies. (a) Ablation studies on proposed modules. (b) Comparisons of rendered normal maps by Gaussian Splatting.
Main Components. Table 4 and Figure 8 (a) present our ablation studies conducted on the Scannet dataset. The first two lines of the table indicate that 2DGS is more suitable for mesh reconstruction tasks compared to 3DGS. Moreover, by incorporating WCD loss and normal prediction network, 2DGS can align better with the actual surfaces. These enhancements significantly improve surface fidelity and reduce geometric artifacts in the final output. Overall, these results demonstrate that each proposed component plays a complementary role in boosting the reconstruction quality.
Gaussian normals. To further demonstrate the necessities of the normal prediction network, we present comparisons on normal maps rendered by Gaussian Splatting in Figure 8 (b). Following [chen2023neusg], we define the normal vector of 3DGS as the direction of the axis which has the minimum value of scales. 2DGS suffer from more significant degradation in normals compared to 3DGS when no regularization is applied. After incorporating the normal prediction network and its corresponding regularization loss, the predicted orientations of 2DGS become more accurate.